A collection of 105 coins consists of 1.00 and 5.00. If the total value is 205.00, Find the number of coins of each denomination in the collection
Answer by Maths68(1474)

Total Number 1:00 coins = x
Total Number of 5.00 coins=105-x
x*1+5(105-x)=205
x+525-5x=205
4x=205-525
4x=320
4x/4=320/4
x=80
Total Number 1:00 coins = x = 80
Total Number of 5.00 coins=105-x = 105-80=25